Can someone help with this problem please:
Solve the matrix equation for X.
A =
2 0 0
-1 2 0
-2 4 1

B =
4
10
11
AX = B
CHOOSE FROM:
a.
0
8
-9

b.
2
6
-9

c.
3
7
-8

d.
3
6
-10
: Can someone help with this problem please:
Solve the matrix equation for X.
A =
2 0 0
-1 2 0
-2 4 1

B =
4
10
11
AX = B
CHOOSE FROM:
a.
0
8
-9

b.
2
6
-9

c.
3
7
-8

d.
3
6
-10

Answer by Fombitz(1789)

AX=B
X=A^(-1)B
Find the inverse of A.
A=(matrix(3,3,<BR>
2 ,0 ,0<BR>
,-1, 2, 0,<BR>
-2, 4, 1  ))
A^(-1)=(1/4)*(matrix(3,3,<BR>
2 ,0 ,0<BR>
,1, 2, 0,<BR>
0, -8,4  ))

Multiply rows of the inverse matrix by the column matrix to get the product.
X[1]= (1/4)*(2(4)+0(10)+0(11))=2
X[2]= (1/4)*(1(4)+2(10)+0(11))=6
X[3]= (1/4)*(0(4)+(-8)(10)+4(11))=-9
.
.
.
(matrix(3,3,<BR>
2 ,0 ,0<BR>
,-1, 2, 0,<BR>
-2, 4, 1  ))*(matrix(3,1, <BR>
2,6,-9 ))=(matrix(3,1, <BR>
4,10,11 ))
.
.
.
b.)
X=(matrix(3,1, <BR>
2,6,-9 ))